OpenStudy (anonymous):

Assume that y varies inversely as x. Find x when y= 6 if y = 4.2 when x = 10. x =

OpenStudy (phi):

Assume that y varies inversely as x that means \[ y = \frac{k}{x} \]where k is some number. to find k, we use the info they gave you: y = 4.2 when x = 10. replace y with 4.2 and x with 10 in the formula: \[ 4.2 = \frac{k}{10} \] "solve" for k, by multiplying both sides by 10: \[ 4.2 \cdot 10= \frac{k}{10}\cdot 10\\ 42=k \] k is 42, so your formula is \[ y = \frac{42}{x} \] Now Find x when y= 6

OpenStudy (jack1):

my bad, i stopped reading after "when y=6"

OpenStudy (phi):

replace y with 6 in the formula \[ 6 = \frac{42}{x} \] multiply both sides by x (to get the x "up top") \[ 6x = \frac{42}{x} \cdot x \\ 6x=42\] now divide both sides by 6 to find x x= 42/6